An apple juice producer buys all his apples from a large orchard. The amount of juice squeezed from each of these apples is approximately normally distributed with a mean of 4.70 ounces and a standard deviation of 0.40 ounce. Suppose a sample of 25 apples is selected, what is the probability that the sample mean will be at least 4.60 ounces?
Answer #1

Solution :

Given that ,

mean = \mu = 4.70

standard deviation = \sigma = 0.40

\sigma\bar x = \sigma / \sqrt n = 0.40 / \sqrt 25 = 0.08

P(\bar x\geq 4.60) = 1 - P(\bar x\leq 4.60)

= 1 - P[(\bar x - \mu \bar x ) / \sigma \bar x\leq(4.60 - 4.70) / 0.08]

= 1 - P(z \leq -1.25)

= 1 - 0.1056

= 0.8944

Probability = 0.8944